OpenStudy (ray10):

Express \[z=8+8i\] in polar form and hence find \[z ^{8}\]? I've done the first part, just don't know how to do the second part

OpenStudy (anonymous):

Using de Moirves theorem thingy

OpenStudy (anonymous):

I think it was something like:\[ [r(\cos(x)+i\sin(x))]^n = r^n(\cos(nx)+i\sin(nx)) \]

OpenStudy (ray10):

but they ask for the answer in polar form and I get a massive number :S can you please help me out with the theorem?

OpenStudy (ray10):

@wio could you show me how the theorem works please? I keep getting a ridiculous number :/

OpenStudy (anonymous):

What do you have for \(x\) or I suppose it's usually \(\theta\)

OpenStudy (anonymous):

What are \(r\) and \(\theta\)?

OpenStudy (ray10):

\[\theta \] comes up as \[\frac{ \Pi }{ 4 }\] r comes up as \[8\sqrt{2}\]

OpenStudy (anonymous):

So what do you get for \(8\theta\) and for \(r^8\)?

OpenStudy (ray10):

\[r ^{8} = 128\] and \[8\theta \] = \[2\Pi \]

OpenStudy (ray10):

it asks for the polar form of; \[z ^{8} \]

OpenStudy (anonymous):

Okay so you should be getting \[ 128(\cos(2\pi)+i\sin(2\pi)) = 128((1)+i(0)) = 128 \]

OpenStudy (ray10):

does that make the polar form; \[128CIS(2\Pi)\]

OpenStudy (anonymous):

The polar form is just \(128\). The imaginary part is \(0\).

OpenStudy (ray10):

and the angle is what then?

OpenStudy (anonymous):

The angle is \(2\pi = 0\).

OpenStudy (anonymous):